Your myopic eye cannot focus on objects that are more than 16.5 cm away or closer than 4.5 cm. When you put on glasses for normal distant vision, your new near point is
Answer #1

lens formula

1/f = 1/p +1/q

1/f = 1/∞ + 1/16.5

f = 16.5 cm

so focal length f = 16.5 cm

again using same lens formula

1/ q = 1/f -1/p

1/q = 1/16.5 -1/4.5

q = -6.18 cm answer

minus sign means image is same side of the lens

****************************************************************************
Goodluck for exam Comment in case any doubt, will reply for sure..